Question 5 (15 pts). – Dave Green, an analyst at Hyvee, shared his forecasts of the daily Tide sales for days 1 to 15 with you as given in the table below (actual sales for days 1 to 15 is already given to you in question 4 above). Using Dave’s forecasts given below answer the following questions:

Day 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15
Dave's Forecast 20 27 14 14 8 11 18 15 8 24 12 26 24 27 24

a) (10 pts): Build a control chart to monitor Dave’s forecast errors for days 1 to 15 (use 3 standard deviation control limits). Plot the forecast errors on the control chart you build.

b) (5 pts): Using the control chart you plot in part (a), discuss whether Dave’s forecast errors are in control or not.
Answer #1

The errors in forecast make the c chart to be appropriate for the case.

average error cbar = 272 /15 = 18.133
UCL of c chart = cbar +3 x (cbar)1/2 = 18.133 + 3 x(18.133)1/2 = 30.907

LCL of c chart = cbar -3 x (cbar)1/2 = 5.35 ( however, a zero value of errors is always desirable)

(b) Since all the observed values are under control limit of 30.907, the process is in control